### v1

# Principles for Open Data Curation: A Case Study with the New York City 311 Service Request Data

Abstract
In the early 21st century, the open data movement began to transform societies and governments by promoting transparency, innovation, and public engagement. The City of New York (NYC) has been at the forefront of this movement since the enactment of the Open Data Law in 2012, creating the NYC Open Data portal. The portal currently hosts 2,700 datasets, serving as a crucial resource for research across various domains, including health, urban development, and transportation. However, the effective use of open data relies heavily on data quality and usability, challenges that remain insufficiently addressed in the literature. This paper examines these challenges via a case study of the NYC 311 Service Request dataset, identifying key issues in data validity, consistency, and curation efficiency. We propose a set of data curation principles, tailored for government-released open data, to address these challenges. Our findings highlight the importance of harmonized field definitions, streamlined storage, and automated quality checks, offering practical guidelines for improving the reliability and utility of open datasets.
1 Introduction
In the early 21st century, the open data movement began to take shape, driven by the fundamental belief that freely accessible data can transform both societies and governments. This movement champions the principles of transparency, innovation, and public engagement. A landmark in this journey was the launch of the United States’ Data.gov portal in 2009 (General Services Administration,, 2024), a pioneering platform in making government data widely accessible. Shortly afterwards, the European Union followed suit, unveiling its Open Data Portal (European Union Plublications Office,, 2024) in 2012, further cementing the movement’s global reach. Furthermore, the World Bank’s Open Data initiative, initiated in 2010, stands out as a comprehensive repository for global development data, available at World Bank Open Data (World Bank,, 2024). These initiatives represent significant strides in democratizing data, removing barriers that once kept valuable information regarding government performance in silos. Their collective impact extends beyond mere data sharing to fostering a culture of openness that benefits individuals, communities, governments, and economies worldwide (Barns,, 2016; Wang and Lo,, 2016). Despite these global efforts, the effective use of open data hinges on addressing critical challenges in data curation, such as ensuring quality, consistency, and usability.
The City of New York (NYC) has emerged as a leader in the open data movement, marked by the enactment of the Open Data Law in 2012 (Zuiderwijk and Janssen,, 2014). This landmark legislation led to the creation of the NYC Open Data Portal (NYC Open Data Team,, 2024), which today hosts an impressive array of 2,700 datasets across 80 different city agencies. This wealth of data serves as a powerful tool for researchers and policymakers, significantly enhancing local government transparency. Popular datasets include information on restaurant health inspection violations, car crashes, high school and college enrollment statistics, jail inmate charges, and the location of City-wide free Internet access points. These datasets have been applied in civil life in various ways, such as mapping car crashes involving pedestrians and visualizing high school and college enrollment trends. Furthermore, they have enabled significant research across multiple domains, including health (Cantor et al.,, 2018; Shankar et al.,, 2021), urban development (Neves et al.,, 2020), and transportation (Gerte et al.,, 2019), aiding in the understanding and addressing of complex urban challenges. For example, NYC’s 311 data has been used to optimize resource allocation in urban planning and improve emergency response times, showcasing its practical relevance in real-time decision-making.
However, the open data presents substantial curation challenges. Data curation, the process of organizing, maintaining, and ensuring the quality of datasets, plays a crucial role in maximizing the utility of open data. Proper curation ensures that datasets remain consistent, accurate, and useful for diverse applications. For example, well-curated data is essential for machine learning systems, which require high-quality data to produce reliable insights (Polyzotis et al.,, 2019; Jain et al.,, 2020). A critical component of data curation is data cleaning, which involves identifying and rectifying inconsistencies, errors, and inaccuracies in datasets. While open data initiatives have made vast amounts of data available, ensuring the reliability and utility of such data hinges on rigorous data cleaning processes. Efficient usage of software tools can significantly streamline this aspect of curation (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018). Poor curation may result in issues such as missing data, formatting errors, or inconsistent values, leading to biased or inaccurate outcomes (Geiger et al.,, 2020). This is particularly critical in domains where machine learning is applied to sensitive tasks, such as public health or policy (Rahm and Do,, 2000).
Research into data curation has explored these challenges in-depth. Among the earliest discussions, Witt et al., (2009) focused on developing data curation profiles tailored to specific contexts, setting a precedent for targeted data management strategies. Addressing broader challenges in data sharing and management, Borgman, (2012) highlighted the complexities of research data distribution, emphasizing the need for robust strategies. This is complemented by Hart et al., (2016), who outlined essential principles for effective data management, particularly emphasizing meticulous curation practices. The utility of curated open data is vividly illustrated in public health and global challenges, where Cantor et al., (2018) demonstrated the utility of curated data in evaluating community health determinants, and Shankar et al., (2021) observed its critical role during the COVID-19 pandemic in managing collective responses. Despite these advances, there remains a notable gap in providing actionable frameworks for managing large-scale municipal datasets.
By addressing this gap, this study transitions from theory to practical insights, aiming to establish principles that enhance both the reliability and utility of open government data. The contributions of this paper are twofold. First, we delve into the specifics of data curation challenges using the NYC 311 Service Request (SR) Data as a case study. This dataset serves as a prime example for examining key issues in data curation, including data validity, consistency, and curation efficiency. We illustrate these points with live examples drawn from our processing of the 311 SR data. Secondly, building upon insights gained from this case study, we propose a set of data curation principles tailored for government-released open data. These principles are designed to address the unique challenges and requirements observed in the curation of such datasets.
The paper is organized as follows. Section 2 provides an overview of the history of the NYC 311 SR system and presents a summary of SR counts over a 2-year period from 2022 to 2023. Section 3 delves into specific data cleansing challenges affecting data quality and curation efficiency, including structural problems, adherence to the data dictionary, and the presence of missing, blank, or N/A entries. This section also investigates data field compliance with reference or acceptable values, highlights logical inconsistencies, and examines some concerning patterns in the data. We explore the balance between precision and accuracy and identify duplicate or redundant fields, along with observations on the Data Dictionary (NYC Open Data Team,, 2023). Section 4 offers practical recommendations for mitigating or resolving these issues, while Section 5 encapsulates key insights and discusses the broader implications of our findings.
2 NYC 311 Service Request (SR) Data
The NYC 311 service, a critical component of New York City’s public engagement and service response framework, serves as a centralized hub for non-emergency inquiries and requests. Introduced in 2003, the NYC 311 system was designed to streamline the city’s response to non-emergency issues, ranging from noise complaints to street maintenance requests. Initially a phone-based call center, the system evolved into a comprehensive data management platform handling millions of requests annually. Key milestones since its launch in 2003 include the addition of online and mobile channels in 2009, a record monthly high of 348,463 SRs in August 2020 (associated with the the COVID-19 pandemic), the 2021 expansion to include the Metropolitan Transit Authority (MTA) NYC subway system (the biggest expansion in 311 history), and a yearly record of 3.23 million SRs in 2023.
Today, the NYC 311 data system manages over 3 million service requests (SRs) per year. This data is publicly accessible through the NYC Open Data Portal, which provides tools for querying, grouping, aggregating, visualizing, and exporting results. The NYC Office of Technology and Innovation (OTI) provides technical support for the 311 application, the open data infrastructure, mobile & web applications, and the public-facing data portal. Input data for the 311 system is sourced from 16 different NYC Agencies. Some of these Agencies utilize the core 311 software, while others transfer data to the 311 Open Data Warehouse from their in-house systems.
The impact of NYC 311 data extends beyond operational efficiency; it has become instrumental in shaping City governance and community engagement. Open data not only ensures governmental transparency but also empowers civic developers, the general public, and policy-makers (Minkoff,, 2016; O’Brien et al.,, 2017; Kontokosta and Hong,, 2021). The data has been pivotal in providing advice on shelters during emergencies, handling inquiries during the COVID-19 pandemic, enforcing standards between landlords and tenants, reallocating taxi routes based on analyses by the Taxi and Limousine Commission (TLC), and improving responsiveness across City Agencies. It also supports a wide range of studies in resource allocation and emergency response strategies (Zha and Veloso,, 2014; Raj et al.,, 2021), social equity in service provision (White and Trump,, 2018; Kontokosta and Hong,, 2021), and urban challenges such as noise pollution (Dove et al.,, 2022) and street flooding (Agonafir et al.,, 2022).
Our investigation uses a 311 SR dataset covering 2-years (2022-2023), and was collected on 15 September 2024. It is to conduct analysis in areas such as data consistency, data validity, redundant fields, etc.. Instructions on downloading this file are available in the supplemental material. Characteristics of this two-year 311 SR dataset are as follows.
  • The dataset is 3.7 GB in size and contains over 6.4 million rows. It exports as a CSV file.
  • There are 41 columns (fields) of data for each row, with each row representing a single SR. To facilitate processing, column names were made lower case and spaces were replaced with underscores, e.g. complaint_type. This naming convention is used throughout this paper.
  • Each SR has four date fields:
    • created_date
    • closed_date
    • resolution_action_updated_date
    • due_date
  • There are two borough fields, borough & park_borough, which appear to be duplicates.
  • There are seven street fields; two pair of which appear to be duplicates:
    • incident_address
    • street_name
    • cross_street_1
    • cross_street_2
    • intersection_street_1
    • intersection_street_2
    • landmark
  • In addtion to incident_address, each SR has eight additional geographic fields:
    • latitude & longitude
    • location
    • street_name
    • landmark
    • block
    • x_coordinate_state_plane & y_coordinate_state_plane.
  • A free-form text field, resolution_description, which supports 930 characters including commas and special characters; often problematic for automated processing.
During the 10-year period from 2014 to 2023, NYC’s 311 system experienced a 50% increase in the number of SRs, reflecting a growing public reliance on and usage of the 311 system. (During this 10-year time frame, the population of NYC grew just 0.6%.). The most significant increase in activity occurred in 2020, when the volume of requests surged to 3.23 million during the COVID-19 pandemic, illustrating the essential role of the 311 system in supporting NYC residents during times of crisis. This steady rise in SRs can be at least partially attributed to the increased accessibility of the 311 system via online and mobile platforms, as well as heightened public awareness of the service. While the spike in 2020 was exceptional, the broader trend indicates sustained growth in SR volume in future years, highlighting the system’s expanding role in managing both routine City operations and extraordinary events.
Figure 1 provides a breakdown of SRs by the responsible Agency, showing the cumulative percentage of SRs handled by each agency over the 2-year time frame. Note that the distribution of SRs is heavily concentrated among a few key agencies. The six largest agencies are:
  1. 1.
    New York Police Department (NYPD): 43% of all SRs
  2. 2.
    Housing Preservation and Development (HPD): 21%
  3. 3.
    New York City Department of Sanitation (DSNY): 10%
  4. 4.
    Department of Transportation (DOT): 7%
  5. 5.
    Department of Environmental Protection (DEP): 5%
  6. 6.
    Department of Parks and Recreation (DPR): 4%
Collectively these six City Agencies handle over 90% of the total SR volume, indicative of the critical role they play in managing public concerns; ranging from noise complaints and housing issues to sanitation and transportation issues. The remaining 10% of SRs are distributed across 10 additional Agencies. This concentration of SRs to these “big six” Agencies underscores the necessity for optimal data curation processes within these high volume Agencies.
Figure 1: 311 SR counts by agency for the time period 2022–2023.
The dataset contains 210 different complaint_types. Figure 2 illustrates how the Top 20 complaint_type(s) account for 70% of all SRs. The distribution of complaints is skewed, with a small number of issues dominating total SR volume. Of special interest are noise-related complaints, of which there are eight different types including vehicle, residential, commercial, and street noise. When combined, noise-related SRs make up 22% of all complaints; the most frequent issue in the NYC 311 system. Other prominent categories include illegal parking, heat/hot water complaints, and blocked driveways. The cumulative percentage curve reveals that the Top 20 complaint_type(s) comprise 70% of SRs, while the remaining 190 other complaint_type(s) are spread thinly across the remaining 30%. This suggests that improving responses to the most frequent complaints could have an outsized impact on overall service efficiency and resident satisfaction. This chart also provides insights into the operational pressures faced by those City Agencies responsible for handling these high-volume complaints.
Figure 2: Top 20 complaint types for the 311 SR during 2022–2023.
3 Data Cleansing Issues
Data cleansing refers to the process of identifying and rectifying errors, inconsistencies, and inaccuracies within datasets to ensure they are of high quality and reliable for analysis (Maletic and Marcus,, 2005; Hosseinzadeh et al.,, 2023). The process typically involves removing duplicate records, handling missing or incomplete data, correcting mislabeled or inaccurate entries, and standardizing data formats (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018). In the context of open data, cleansing is especially important as open datasets often come from diverse, often uncoordinated sources, leading to variations in data quality, completeness, and consistency. Without thorough data cleansing, the utility of open data can be limited and perhaps untrustworthy, affecting its reliability for research, policy making, and innovation. The main purpose of cleansing open data is to ensure that it is accurate, consistent, and usable a cross multiple analytical platforms and by various stakeholders with a myriad of purposes. Data cleansing improves the trustworthiness of the data and enables more accurate analysis, better decision making, and the improved integration of data into machine learning models or other systems.
Many quality criteria are employed to ensure high-quality data processing. One of the primary efforts is data validation, which includes several critical checks. For instance, mandatory fields must not be left empty, ensuring that key information is always captured. Additionally, certain fields must conform to specific data types, such as numeric, character, or date formats; typically outlined in a Data Dictionary. An essential aspect of validation is domain compliance, where fields must adhere to a predefined set of values such as statuses, state names, or zip codes. Structural errors can also play a significant role in data quality, particularly when naming conventions or data entries are inconsistent. Common issues include fields that do not appear in the Data Dictionary and inconsistent usage of data fields. Furthermore, redundant (or irrelevant) fields can clutter datasets, reducing efficiency, creating confusion, and introducing errors. Logical inconsistencies are another consideration, such as related fields that violate expected relationships, e.g. a “due date” that precedes the “created date.” Lastly, the balance between accuracy and precision is crucial, as both must be managed to ensure meaningful data.
We identify the presence of issues in the NYC 311 SR open dataset; not attempting to solve them, but rather to highlight their scope, magnitude, and potential impact. We would recommend that a corrective solution should be undertaken only after further investigation as to the why and how such issues might come about, including detailed discussions with the originating agency as to whether or not the issue is an actual error, correct, or holds some other status.
3.1 Structural Issues
Structural issues refers to how data is organized, formatted, or presented within a dataset. Structural issues can cause difficulty in analyzing the data effectively, often necessitating extensive data cleaning. We identified three primary structural issues in the NYC 311 dataset: undocumented fields, inconsistent representations of missing data, and daylight saving time anomalies.
Undocumented Fields:
The 311 SR The 311 SR Data Dictionary identifies 41 data columns (fields) and provides related information for each. However, there are six additional fields that can be optionally downloaded if manually selected. These fields are visible in the portal’s “Column manager” widget, labeled with the prefix “@computed_region.” and include:
  • zip_codes
  • community_districts
  • borough_boundaries
  • city_council_districts
  • police_precincts
  • police_precinct
While one can likely infer the meaning of these computed fields, uncertainty remains as to their derivation and usage. These six fields are not included in the 311 SR Data Dictionary. Additionally, our analysis revealed that these six “computed” fields have significant data validity issues, leading us to categorize these six fields as experimental and not for public use. An update to the Data Dictionary would seem appropriate. Accordingly, we omit these fields in our analysis.
Representation of Missing data:
We found that the 311 SR dataset exhibited inconsistent approaches when presenting missing data. We discovered different representations of missing data: nulls, spaces, “NA”, “N/A”, and “<NA>”. This variety of representations complicates programming efforts and increases the likelihood of introducing errors in data analysis (typically when selecting data for blank, NA, or missing fields). A more consistent approach to treating missing data should be employed to address this issue.
Challenges with Daylight Saving Time (DST):
The 311 SR dataset captures local NYC times. Accordingly, twice each year there are changes to the local time as a result of observing DST. This can introduce some unexpected results. For example, on both March 13, 2022 and March 12, 2023 there are no times between 1:59AM and 2:59AM. The 2:00AM hour on those dates is simply missing owing to the clocks moving forward; at 2:00AM it instantly becomes 3:00AM. This creates a gap in the hourly distribution of SRs on those days, and exaggerates the lifespan of SRs, albeit by a single hour. Similarly on both November 6, 2022 and November 5, 2023, the hour of 1:00AM - 2:00AM is repeated twice. This can result in a bizarre scenario where, for example, an SR can be created at 1:59AM and then be closed 30 minutes later at 1:29AM; seemingly nonsensical. Utilizing UTC time for date-time stamps would help address this issue.
3.2 Missing Data by Field
Understanding the absence of data by field greatly aids analysis. For example, determining if SRs were closed before their due_date, would be challenging as 99.6% of the due_date entries are blank as only DSNY appears to use this field. Blank (or N/A values) divide into three groups:
  • Mostly Empty: 93-99.9% blank;
  • Partially Empty: 40-4% blank;
  • Few/None Empty: 2-0% blank.
The Mostly Empty category includes: taxi_company_borough, due_date, road_ramp, and bridge_highway_name. The Partially Empty category includes: location_type, landmark and cross_street_1 & _2. The Few/None Empty category includes created_date, complaint_type, agency, and status. Note that several of the Mostly Empty fields are used by only one or two Agencies (See: Table 1).
Table 1: Fields used by only one or two Agencies with limited non-blank entires
Field # of entries % non-blank Agencies using
taxi_company_borough 3,567 0.06% TLC
vehicle_type 18,642 0.29% TLC, NYPD
due_date 24,414 0.38% DSNY
taxi_pick_up_location 64,954 1.02% TLC
Figure 3: Number and percentage of empty/blank entries for the 311 SR data during 2022–2023.
For some analysis efforts, it may be prudent to inquire as to why some fields are mostly/almost always blank. Figure 3 is a graphic depiction of the percentage of empty (blank & N/As) for each field. It illustrates the grouping into Mostly Empty, Partially Empty, and Few/None Empty segments.
Table 2 highlights the varied usage of certain fields across NYC agencies, revealing both specialization and redundancy. For instance, landmark and facility_type are widely used, indicating their importance in classifying SRs. In contrast, fields like taxi_company_borough and vehicle_type are used exclusively by TLC, suggesting that such fields, which serve a narrow purpose, could be better stored in separate, agency-specific tables. This approach would reduce the size and complexity of the primary dataset, making it easier to maintain and analyze while preserving the ability to perform specialized analyses when necessary. Streamlining through harmonization and validation would further enhance consistency and optimize cross-agency analyses.
Table 2: Sample field usage by select agencies among the 6.4 million SRs in 2022–2023.
Field DEP DHS DOB DOT DPR DSNY TLC
landmark 2935 70185 282 135150 254218 424284 63133
facility_type 175549 0 51575 159667 0 66012 0
due_date 0 0 0 0 0 32097 0
vehicle_type 0 0 0 0 0 0 1696
taxi_company_borough 0 0 0 0 0 0 4025
taxi_pick_up_location 0 0 0 0 0 0 72558
bridge_highway_name 15 21012 0 7451 0 613 977
road_ramp 15 2588 0 7450 0 613 999
3.3 Validating Data for Acceptable Values
Any analytic effort must ensure that fields containing invalid values are identified and isolated from the analysis. For example, the latitude and longitude fields were found to all fall within the geographic boundaries of New York City. Likewise, the unique_key field was indeed unique, as required. Unfortunately, the 311 SR Data Dictionary specifies very few domains of acceptable values for other fields. Nonetheless, the following fields were tested and found to comply with their expected domains, as determined by public usage and as evidenced in other historical datasets:
  • address_type
  • status
  • borough
  • borough_boundaries
  • park_borough
  • data_channel
  • vehicle_type
  • city_council_district
However, some fields lacked compliance to a domain of allowable values, as outlined next.
Zip Codes:
All incident_zip zip codes should validate against the USPS database, which contains 44,173 valid zip codes. We discovered that the incident_zip field has 0.07% invalid entries, resulting in 4,374 errors. This includes zip codes such as “12345”, “10000”, and “99999”. If you group the invalid zip code entries by percentage by agency, it mirrors the overall breakdown of SRs by Agency, potentially indicating a systemic problem.
Created and Closed Dates:
With the created_date and closed_date fields, one might expect that these fields would be automatically populated by the application software. Thus when creating or closing an SR the the software would automatically populate these fields. Unfortunately, this does not seem to be the case, as several anomalies exist in the created_date and closed_date fields that an automated process would normally exclude, including:
  • 12,450 SRs with a closed_date before created_date creating a nonsensical “negative duration”.
  • Eight SRs with created_date(s) or closed_date(s) in the distant past, specifically 1900-01-01. Despite few of these entries, such extreme values can greatly distort analytical efforts.
  • 163,720 SRs with identical created_date(s) and closed_date(s) to the second, creating a nonsensical “zero duration”.
  • 247,000 SRs closed and/or created exactly at midnight, to the second.
Closed before Created (Negative Duration):
Citizens, NYC Government Officials, and Agencies use the created_date(s) and closed_date(s) to measure the duration of SRs, a useful surrogate to measure an agency’s responsiveness. While duration is not directly present in the dataset, it is easily computed as the difference between closed_date and created_date. Duration is one of most frequently used 311 performance metrics and is frequently analyzed, such as determining if one NYC borough receives faster Agency responses than another borough for certain complaints. There are 12,450 SRs where the closed_date precedes the created_date, generating nonsensical “negative durations.” Although this represents only 0.2% of the dataset, these errors can significantly impact response time analyses. Eight SRs with extremely large negative durations, all originating from the Department of Homeless Services (DHS); all contain an entry of 1900-01-01 as the closed_date. This results in very large negative durations exceeding --44,601 days (or 122 years). Many other SRs have negative durations in excess of --300+ days. Such anomalies, though rare, can significantly skew statistical results if not addressed during data cleansing. Accordingly, these SR rows are removed from our analysis.
Excluding the extreme negative values (any that are lower than --730 days), Figure 4(a) shows the broad spread of negative duration SRs. While there are few outliers, the magnitude of these negative durations (some lower than --300 days) is troubling and can produce bizarre analytical results such as greatly distorting the mean response time for a specific complaint category.
(a) Distribution of SR durations.
(b) Post-closed resolution_action_update_date(s) ¿30 days
Figure 4: Violin plots for negative SR durations and post-closed resolution updates (2022–2023.
Table 3 summarizes the negative durations of the largest magnitude (in days). The data suggests that the negative duration issue is predominantly a problem within the Department of Transportation (DOT), where 95% of these errors occur.
Table 3: Sample negative durations [excluding extreme values that are lower than --730 days]
created_date closed_date Duration (days) Agency
2023-01-27 14:40:00 2022-01-14 14:40:00 --378 DOT
2023-01-18 10:06:00 2022-01-12 10:06:00 --371 DOT
2023-01-27 14:36:00 2022-01-22 14:35:00 --370 DOT
2023-01-11 11:10:00 2022-01-09 11:10:00 --367 DOT
2023-12-18 03:13:00 2023-01-16 13:10:00 --335 DOT
Identical Created & Closed Dates (Zero Durations):
A more prevalent issue occurs when the closed_date and created_date have exactly the same date-time stamp, to the second. This creates a zero duration for the SR, which is nonsensical. There are 163,720 such zero duration SRs, representing 2.6% of all non-blank data. Additionally, 99% percent of these zero duration SRs occur within five agencies: Department of Health & Mental Hygiene (DHMH), DOT, Department of Buildings (DOB), DSNY, and DEP. This distribution does not mirror the overall SR distribution by agency, suggesting an Agency-specific issue.
Created or Closed at Midnight:
[…]
(a) SRs created exactly on the hour
(b) SRs closed exactly on the hour
Figure 5: 311 SRs Created & Closed exactly on the hour (hh:00:00) for 2022–2023.
These unusual patterns of SR creation/closure at midnight suggest the presence of a bulk create/close software process, perhaps one that automatically assigns time stamps of midnight (00:00:00) to batches of SRs prior to sending the data to the 311 database. This behavior distorts the calculated duration of these SRs as well as overall system usage by providing inaccurate dates. Further analysis of the distribution by Agency shows that over 92% of the “closed-exactly-at-midnight” SRs come from just two Agencies: DOB (69%) and DSNY (23%).
Resolution Action Update Date
A similar issue occurs when an SR is modified (and the resolution_action_update_date is altered) well after the closed_date. When an SR is updated, the 311 software automatically populates the resolution_action_update_date. While it is certainly possible, and even routine, to update an SR after it is closed, some of these updates seem to be well beyond any normal or expected timeframe, with some “updates” occurring over 2 years later. As shown in Figure 4(b), there are 10,265 SRs that were updated $>30$ days after the closed_date (Not that Figure 4(b) excludes values $>730$ days, an arbitrary cutoff to exclude infeasible dates such as 1900-01-01. The chart highlights the distribution of these post-closure updates, with a notable concentration of updates occurring within the 30–200 day range. This pattern raises questions about whether such delayed updates are standard operational practice or indicative of a potential issue which warrants further investigation. It should be noted that 84% of these late updates are associated with the TLC agency and 12% with DSNY.
3.4 Accuracy and Precision
The question of precision vs. accuracy arises with the latitude and longitude fields. as well as with the duplicative location field. These fields are expressed as a 14-decimal number, e.g. latitude 40.86769186022511. Given that 1 degree of latitude at the equator is equal to 111.04 kilometers, the “1” at the end of that decimal number represents approximately 1.1 nanometers (1/1,000,000,000 of a meter). (For reference a DNA molecule is approximately 2 nanometers in width.) This representation appears to be a classic case of extreme precision, albeit with limited accuracy. All the latitude, longitude, and location fields display this characteristic.
3.5 Redundant Fields
This analysis revealed several redundant fields, indicting a potential need for consolidation.
latitude/longitude & location:
The location field is a pure concatenation of the latitude and longitude fields albeit with a comma and parentheses added. Arguably, this makes the location field more difficult to use than the individual fields. For example:
40.768456429488--73.9575661888774(40.768456429488, --73.9575661888774)latitude & longitude fieldslocation field
borough & park_borough:
These two fields are fully redundant, as they are 100% matches.
borough & taxi_company_borough:
Despite their names, these fields are almost entirely different, with only a 0.05% match. The taxi_company_borough field is used exclusively by the TLC, indicating a need for consultation with that agency to understand the differing uses. It is likely that that the usage of taxi_company_borough is quite different than that of borough.
agency & agency_name:
The agency field contains the abbreviations of City agencies (e.g., NYPD, DOT, DSNY), while agency_name contains the full names. Including both seems redundant, especially given the space taken up by full names in the CSV file and the fact that the agency abbreviations are easily recognized and well understood.
landmark & street_name:
Per the 311 SR Data Dictionary, the landmark field is intended to refer to any noteworthy location, including but not limited to parks, hospitals, airports, sports facilities, and performance spaces. However, most entries are actually street names, with a 62% match to the street_name field. Many of the non-matches (excluding blanks) also appear to be logical matches, differing only in spelling or formatting (e.g., “NINTH AVE” vs. “9 AVE”). This indicates a possible misuse of the landmark field and highlights significant redundancy.
cross_street_1 & 2 & intersection_street_1 & 2:
These two street pairs are used to help identify the incident address. We found 88% of both of these two street pairs (1 & 2) to be duplicates. Such a high level of duplication raises the question of which field-pair should be trusted; and if these street pairs are used differently, how should they be treated? It is possible that these two street pairs originated from different agencies that use different variable names to represent the same information, but no documentation exists to clarify this relationship. Nonetheless, these two street-pairs are highly redundant.
We attempted to standardize street names to reconcile entries that were clearly identical but differed due to spelling or naming conventions. For example, we sought to match entries such as KANE PL and KANE PLACE or AVE P and AVENUE P. However, this standardization effort was ultimately abandoned, as it resulted in less than a 1% increase in matched street names.
3.6 Reducing File Size
By removing duplicate and near-duplicate fields, one can reduce file size by 39.5%, equating to a reduction of 1.4GB, a significant amount. A smaller file size means faster downloads, reduced storage impact, and often simpler data analysis. Below is a list of duplicate and near-duplicate fields which we believe could be removed from the dataset with minimal data loss.
  • agency_name: Since each row contains the agency field, which uses clear and well understood abbreviations, we believe the agency_name can be removed with minimal data loss.
  • park_borough: This field is a 100% match with the borough field. It can be deleted without data loss.
  • location: This field is a straight concatenation of the latitude and longitude fields. It can be deleted without data loss.
  • cross_street_1/2 and intersection_street_1/2: These two street pairs exhibit an 88% match. Although we cannot confirm the accuracy of either of these pairs, we recommend deleting the intersection_street_1/2 fields, acknowledging the concomitant data loss.
Additionally, certain fields may not be useful for broad analysis due to their sparse data population. This applies to the fields below which are 99% (or greater) blank.
  • taxi_company_borough
  • road_ramp
  • vehicle_type
  • due_date
  • bridge_highway_direction
  • bridge_highway_name
  • bridge_highway_segment
  • taxi_pick_up_location
While such sparsity may limit their use in broad analyses, these fields can still contain valuable insights for specific agency-level studies, given that even a 1% non-blank rate equates to approximately 64,000 rows in this 6.4 million-row dataset. Proper handling of sparse fields, such as storing them in separate tables or using them in targeted analyses, could improve data accessibility and efficiency without compromising the dataset’s broader utility.
Field Encoding:
Further size reduction is possible by encoding selected categorical variables Using complaint_type as an example, instead of repeatedly storing the full text of each possible value, one could use a numeric representation. For example, “NOISE - STREET/SIDEWALK” could be represented by a numeric code. For this dataset this code occurs over 300,000 times. Encoding not only reduces file size, but would likely improve processing efficiency by using numeric values instead of text strings. Moreover, standardized categorical inputs often prevents misspellings and invalid entries, ensuring cleaner, more reliable data. This approach is particularly useful in large datasets, where frequently repeated text-based categorical variables significantly increases storage requirements.
One promising alternative to reducing data storage for this large 311 SR dataset (3.7GB) would be employment of the cross-platform Apache Arrow package. Apache Arrow (arrow.apache.org) is an innovative, cross-language development platform designed for in-memory data. It defines a standard for representing columnar data, enabling efficient analytics and real-time data processing. As such, it offers dramatic improvements in file storage size, as well as efficient analysis and cross-platform support for R, Python, and Julia programming languages. Remarkable storage size reduction (within several orders of magnitude) is possible (Bates and Yan,, 2024).
3.7 Data Dictionary
The 311 SR Data Dictionary needs updating. We found several discrepancies between the Data Dictionary and the actual data. Such errors could potentially mislead analytic efforts. Examples::
Data typing:
While the NYC Open Data portal displays the data types of various fields (aka “columns”), the Data Dictionary does not. For example, in the Data Dictionary, the incident_zip field is specified as “text.” This may not be the most suitable data definition. Perhaps a better dictionary description would be “categorical” with a note indicating that the data must contain only numeric values and is not subjected to arithmetic operations.
Domain of Legal Values:
Many of the data fields have domains of legal values that are missing, incomplete, or inaccurate. For example, the Data Dictionary indicates that the status field takes on values of assigned, canceled, closed, or pending. However, we observed additional values such as in progress, started, and unspecified. Additionally, no SRs were found with a status of canceled. Similarly, the Data Dictionary for the address_type field lists as acceptable values address, blockface, intersection, latlong, and placename. However, we found additional values such as bbl and unrecognized. Additionally, no latlong values were observed. Other fields, such as facility_type, vehicle_type, taxi_pick_up_location, textttroad_ramp, and city, exhibit similar inaccuracies in their specified domain values.
Missing data fields:
As mentioned earlier, six @computed fields that are present in the data but not included in the Data Dictionary need to be documented and perhaps labeled as “experimental” or “not for official use”.
4 Government Open Data Curation Recommendations
Building upon insights from our case study of the NYC 311 SR data, we propose a set of data curation principles tailored for government open datasets. These principles are designed to address the unique challenges observed in the curation of such datasets, ensuring they remain reliable, consistent, and useful for public services, research, and other applications. To facilitate implementation, the principles are prioritized based on their impact and feasibility, offering a clear roadmap for improving data curation practices.
4.1 Principles
Principle 1: Establishing Cross-Organizational Consistency
A key challenge in government released datasets is the harmonization of data fields across various Agencies, and maintaining that consistency over time. Inconsistent field definitions, data formats, or value domains can lead to erroneous analyses and hinder cross-agency collaboration. To address this, Agencies should adopt standardized naming conventions, value domains, usages, and formats across all datasets. Regular audits and data harmonization processes should be implemented to ensure consistency across long-term datasets, particularly as Agency structures and reporting standards evolve. For instance, ensuring that historical datasets reflect the same field naming conventions over a span of years can prevent inconsistent analyses. These standards should be promulgated by the Open Data governing authority, following agreements by the relevant parties.
Principle 2: Ensuring Data Entry Accuracy
Data accuracy and validity are fundamental to the utility of open datasets. In the NYC 311 data, issues such as invalid zip codes and nonsensical date entries highlight the need for rigorous validation processes. Governments should establish clear protocols for data entry, ensuring that fields are populated with valid, legally acceptable values. Accurate data entry reduces the risk of introducing systemic errors, particularly in high-volume datasets. To the maximum extent possible,the software should assist in ensuring data entries are accurate and valid. Some examples include:
  • Certain date fields should be promulgated by the application software versus manual entry. Dates should be populated based upon an action in the system, such as a change in status.
  • Select fields should be subjected to automated checks that prevent logical errors, such as a closing date occurring before the creation date.
  • A number of fields have clearly defined domains of legal values. These fields should be validated against a reference dataset upon entry.
Principle 3: Optimizing Storage Efficiency
Efficient storage and data representation are crucial for handling large government datasets. Our research found several duplicate and near-duplicate fields in the 311 SR dataset. Removing these redundancies could reduce the dataset size by 39.5%, improving performance without compromising analytical value. Eliminating redundant fields and encoding categorical variables can significantly reduce file sizes and improve the efficiency of data analysis. Similarly, categorical variables could be encoded in a standardized format to optimize storage and streamline analyses. Reducing file size not only saves storage space but also enhances the performance of queries, downloads, and analysis processes.
Principle 4: Updating Data Dictionaries
Clear and accurate data documentation is essential for the effective use of government datasets. Our study found several instances where the Data Dictionary did not accurately reflect the actual data. For example, undocumented fields and outdated domain definitions in the 311 SR dataset introduced ambiguity, complicating analysis. Domains of legal/allowable values should be referenced in the Data Dictionary. Regular reviews and updates should occur whenever there are changes in data structure, data usage, or content. This practice enhances the ability of external users, including researchers and developers, to work effectively with the data.
Principle 5: Automating Quality Assurance Processes
Automating data validation and quality assurance processes can significantly improve the reliability of government datasets. Certain unusual trends can be identified, and irregularities highlighted. In the NYC 311 case, issues such as large spikes in SR creation and closure at exactly midnight (or noon) likely stem from a cron script process that distorts the accurate capture of SR durations. Real-time validation systems could prevent such errors, ensuring data integrity at the point of entry. By implementing real-time and event-driven validation tools, governments can both prevent such errors from entering the dataset in the first place, as well as highlight anomalous trends. Automated systems can be designed to flag inconsistencies, incorrect values, illogical sequences, or illogical data entries. Automation can also facilitate faster detection and correction of errors, reducing the need for manual intervention and ensuring the data remains clean and reliable as it is continuously updated.
Principle 6: Addressing Data Transparency
Transparency is a core tenet of open data, and governments must ensure that datasets are not only accurate but also easily accessible and readily understandable. This involves providing clear metadata and making datasets available in user-friendly formats. This includes publishing clear metadata, which enables users to understand the dataset’s structure, limitations, and intended use cases. Additionally, governments should communicate transparently about data quality issues, such as notifying the public when errors in SR fields (e.g., erroneous closed_ date(s) are discovered, including an explanation about how such issues will be resolved. By ensuring transparency and keeping the public informed, governments foster trust and promote wider use of datasets for research, civic engagement, and innovation.
4.2 Actionable Steps
While these principles are broad, actionable steps are required to implement them effectively. Prioritizing these steps will help governments address the most critical data curation challenges first. Governments can begin by investing in real-time validation systems, regular updates to data dictionaries, and tools for encoding and standardizing data fields. This would include systems that automate enforcement of data validation rules (e.g., flagging incorrect zip codes), significantly improving data quality. Automated systems can proactively detect and correct common issues, reducing the manual workload. Collaboration with city agencies and stakeholders is crucial to ensure that data standards, formats, and metadata meet the needs of diverse users, including policymakers, researchers, and the general public. Regular communication between agencies and external users can help identify recurring issues and areas where improvements are needed. By facilitating regular communication and collaboration with various stakeholders, governments can ensure data remains relevant and actionable for public use.
Automated Quality Assurance techniques should be employed to monitor the quality of the data, as measured by agreed upon standards, such as the percentage of invalid zip codes. Such tools could capture and display the quality of the data sets as presented on appropriate dashboards. These dashboards provide a real-time snapshot of data quality, making deviations immediately visible and actionable. Deviations in quality would be automatically highlighted, alerting appropriate corrective actions.
Engagement with data scientists and statisticians is key to improving data curation strategies. These experts can help design and refine curation processes, such as creating algorithms for automatic data cleaning, identifying inconsistent data entries, and ensuring the dataset(s) structure is maintained over time. Their involvement also helps to establish best practices for validating incoming data and maintaining consistency across data sources. Feedback, such as the issues reported in this paper, benefit from a clear and accessible channel to the 311 SR data team and other relevant City Agencies. Establishing such feedback loops and other opportunities for public engagement, ensures that recommendations for improving data curation and validation are not only heard and discussed, but implemented. An iterative feedback system enables timely resolution of issues and fosters long-term improvements in data management.
An especially valuable engagement avenue is through events such as NYC Open Data Week, which foster collaboration and innovation across sectors. Our involvement in this project stemmed from such an event, where we were inspired by the potential of open data to enhance public services. Governments should continue these types of initiatives, which not only promote data literacy, but encourage community-driven improvements to datasets. These events can strengthen public engagement and demonstrate the societal value of open data initiative By leveraging the expertise of data scientists, statisticians, stakeholders, and civic organizations, governments can ensure their open datasets remain accurate, transparent, and useful for a wide range of applications.
5 Discussion
Our study highlights the critical role of robust data curation in ensuring the reliability and utility of open datasets, as demonstrated through the examination of the NYC 311 data. As these datasets are applied across diverse fields such as public services and academic research, inconsistencies, missing values, and formatting errors can significantly undermine the insights derived from them. Trustworthy machine learning systems, which rely heavily on high quality data, are especially vulnerable to such issues. Errors in the underlying data, such as biases or inconsistencies, can compromise both the accuracy and fairness of predictive models, impacting real-world decisions in urban planning and policy (Rahm and Do,, 2000; Geiger et al.,, 2020).
While the findings here are specific to NYC 311 data, the principles proposed have broader applicability. Datasets in healthcare, transportation, and environmental monitoring can similarly benefit from these curation strategies, enabling improved outcomes across domains. This research further emphasizes the necessity of harmonizing data across City agencies, particularly resolving discrepancies in field names, formats, usage, and definitions. Consistency in data is essential for trustworthy machine learning, where inconsistencies across sources can lead to distorted outcomes. This challenge is even more pronounced in datasets that span long time periods, as changes in agency structures or reporting standards which can introduce subtle biases and data distortions. Long-term data consistency is crucial for longitudinal studies and predictive modeling, as even minor data shifts can lead to significant deviations in model outcomes, as noted by Rahm and Do, (2000) andBorgman, (2012).
Finally, the intersection of data curation and machine learning for public policy applications opens new avenues for improving governance. High quality data enables machine learning models to better predict trends, allocate resources, and address service delivery issues, such as delays in responding to 311 complaints. The COVID-19 pandemic underscored the importance of real-time, trustworthy data in crisis management, where the accuracy and timeliness of data-driven insights were critical for public health responses. Without proper data curation, machine learning models used during the pandemic would have been compromised, affecting decisions on resource allocation and service provision (Worby and Chang,, 2020; Khemasuwan and Colt,, 2021). Future efforts could focus on automating data curation processes, particularly in real-time data pipelines, to ensure that the data used in machine learning models remains accurate, clean, and reliable (Chu et al.,, 2016; Hurbean et al.,, 2021).
Supplementary Material
The NYC 311 R code used to produce these results, along with additional analytical findings, are available at https://github.com/jun-yan/nyc311clean. Reference data files (311 SR data of 2022–2023, USPS zip codes) are available at https://doi.org/10.6084/m9.figshare.c.7617311.

### v2

@@
 Our findings highlight the importance of harmonized field definitions, streamlined storage, and automated quality checks, offering practical guidelines for improving the reliability and utility of open datasets.
+Data cleansing,
+Data science,
+NYC Open Data,
+Smart city,
+Transparency,
+doi:
+$0$
+keywords:
+volume: 0issue: 0articletype: research-article
 1 Introduction
@@
 This movement champions the principles of transparency, innovation, and public engagement.
-A landmark in this journey was the launch of the United States’ Data.gov portal in 2009 (General Services Administration,, 2024), a pioneering platform in making government data widely accessible.
-Shortly afterwards, the European Union followed suit, unveiling its Open Data Portal (European Union Plublications Office,, 2024) in 2012, further cementing the movement’s global reach.
-Furthermore, the World Bank’s Open Data initiative, initiated in 2010, stands out as a comprehensive repository for global development data, available at World Bank Open Data (World Bank,, 2024).
+A landmark in this journey was the launch of the United States’ Data.gov portal in 2009 (General Services Administration, 2024), a pioneering platform in making government data widely accessible.
+Shortly afterwards, the European Union followed suit, unveiling its Open Data Portal (European Union Plublications Office, 2024) in 2012, further cementing the movement’s global reach.
+Furthermore, the World Bank’s Open Data initiative, initiated in 2010, stands out as a comprehensive repository for global development data, available at World Bank Open Data (World Bank, 2024).
 These initiatives represent significant strides in democratizing data, removing barriers that once kept valuable information regarding government performance in silos.
-Their collective impact extends beyond mere data sharing to fostering a culture of openness that benefits individuals, communities, governments, and economies worldwide (Barns,, 2016; Wang and Lo,, 2016).
+Their collective impact extends beyond mere data sharing to fostering a culture of openness that benefits individuals, communities, governments, and economies worldwide (Barns, 2016; Wang and Lo, 2016).
 Despite these global efforts, the effective use of open data hinges on addressing critical challenges in data curation, such as ensuring quality, consistency, and usability.
-The City of New York (NYC) has emerged as a leader in the open data movement, marked by the enactment of the Open Data Law in 2012 (Zuiderwijk and Janssen,, 2014).
-This landmark legislation led to the creation of the NYC Open Data Portal (NYC Open Data Team,, 2024), which today hosts an impressive array of 2,700 datasets across 80 different city agencies.
+The City of New York (NYC) has emerged as a leader in the open data movement, marked by the enactment of the Open Data Law in 2012 (Zuiderwijk and Janssen, 2014).
+This landmark legislation led to the creation of the NYC Open Data Portal (NYC Open Data Team, 2024), which today hosts an impressive array of 2,700 datasets across 80 different city agencies.
 This wealth of data serves as a powerful tool for researchers and policymakers, significantly enhancing local government transparency.
@@
 These datasets have been applied in civil life in various ways, such as mapping car crashes involving pedestrians and visualizing high school and college enrollment trends.
-Furthermore, they have enabled significant research across multiple domains, including health (Cantor et al.,, 2018; Shankar et al.,, 2021), urban development (Neves et al.,, 2020), and transportation (Gerte et al.,, 2019), aiding in the understanding and addressing of complex urban challenges.
+Furthermore, they have enabled significant research across multiple domains, including health (Cantor et al., 2018; Shankar et al., 2021), urban development (Neves et al., 2020), and transportation (Gerte et al., 2019), aiding in the understanding and addressing of complex urban challenges.
 For example, NYC’s 311 data has been used to optimize resource allocation in urban planning and improve emergency response times, showcasing its practical relevance in real-time decision-making.
@@
 Proper curation ensures that datasets remain consistent, accurate, and useful for diverse applications.
-For example, well-curated data is essential for machine learning systems, which require high-quality data to produce reliable insights (Polyzotis et al.,, 2019; Jain et al.,, 2020).
+For example, well-curated data is essential for machine learning systems, which require high-quality data to produce reliable insights (Polyzotis et al., 2019; Jain et al., 2020).
 A critical component of data curation is data cleaning, which involves identifying and rectifying inconsistencies, errors, and inaccuracies in datasets.
 While open data initiatives have made vast amounts of data available, ensuring the reliability and utility of such data hinges on rigorous data cleaning processes.
-Efficient usage of software tools can significantly streamline this aspect of curation (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018).
-Poor curation may result in issues such as missing data, formatting errors, or inconsistent values, leading to biased or inaccurate outcomes (Geiger et al.,, 2020).
-This is particularly critical in domains where machine learning is applied to sensitive tasks, such as public health or policy (Rahm and Do,, 2000).
+Efficient usage of software tools can significantly streamline this aspect of curation (e.g., Cody, 2017; Van der Loo and De Jonge, 2018).
+Poor curation may result in issues such as missing data, formatting errors, or inconsistent values, leading to biased or inaccurate outcomes (Geiger et al., 2020).
+This is particularly critical in domains where machine learning is applied to sensitive tasks, such as public health or policy (Rahm and Do, 2000).
 Research into data curation has explored these challenges in-depth.
-Among the earliest discussions, Witt et al., (2009) focused on developing data curation profiles tailored to specific contexts, setting a precedent for targeted data management strategies.
-Addressing broader challenges in data sharing and management, Borgman, (2012) highlighted the complexities of research data distribution, emphasizing the need for robust strategies.
-This is complemented by Hart et al., (2016), who outlined essential principles for effective data management, particularly emphasizing meticulous curation practices.
-The utility of curated open data is vividly illustrated in public health and global challenges, where Cantor et al., (2018) demonstrated the utility of curated data in evaluating community health determinants, and Shankar et al., (2021) observed its critical role during the COVID-19 pandemic in managing collective responses.
+Among the earliest discussions, Witt et al.
+(2009) focused on developing data curation profiles tailored to specific contexts, setting a precedent for targeted data management strategies.
+Addressing broader challenges in data sharing and management, Borgman (2012) highlighted the complexities of research data distribution, emphasizing the need for robust strategies.
+This is complemented by Hart et al.
+(2016), who outlined essential principles for effective data management, particularly emphasizing meticulous curation practices.
+The utility of curated open data is vividly illustrated in public health and global challenges, where Cantor et al.
+(2018) demonstrated the utility of curated data in evaluating community health determinants, and Shankar et al.
+(2021) observed its critical role during the COVID-19 pandemic in managing collective responses.
 Despite these advances, there remains a notable gap in providing actionable frameworks for managing large-scale municipal datasets.
@@
 This section also investigates data field compliance with reference or acceptable values, highlights logical inconsistencies, and examines some concerning patterns in the data.
-We explore the balance between precision and accuracy and identify duplicate or redundant fields, along with observations on the Data Dictionary (NYC Open Data Team,, 2023).
+We explore the balance between precision and accuracy and identify duplicate or redundant fields, along with observations on the Data Dictionary (NYC Open Data Team, 2023).
 Section 4 offers practical recommendations for mitigating or resolving these issues, while Section 5 encapsulates key insights and discusses the broader implications of our findings.
@@
 The impact of NYC 311 data extends beyond operational efficiency; it has become instrumental in shaping City governance and community engagement.
-Open data not only ensures governmental transparency but also empowers civic developers, the general public, and policy-makers (Minkoff,, 2016; O’Brien et al.,, 2017; Kontokosta and Hong,, 2021).
+Open data not only ensures governmental transparency but also empowers civic developers, the general public, and policy-makers (Minkoff, 2016; O’Brien et al., 2017; Kontokosta and Hong, 2021).
 The data has been pivotal in providing advice on shelters during emergencies, handling inquiries during the COVID-19 pandemic, enforcing standards between landlords and tenants, reallocating taxi routes based on analyses by the Taxi and Limousine Commission (TLC), and improving responsiveness across City Agencies.
-It also supports a wide range of studies in resource allocation and emergency response strategies (Zha and Veloso,, 2014; Raj et al.,, 2021), social equity in service provision (White and Trump,, 2018; Kontokosta and Hong,, 2021), and urban challenges such as noise pollution (Dove et al.,, 2022) and street flooding (Agonafir et al.,, 2022).
+It also supports a wide range of studies in resource allocation and emergency response strategies (Zha and Veloso, 2014; Raj et al., 2021), social equity in service provision (White and Trump, 2018; Kontokosta and Hong, 2021), and urban challenges such as noise pollution (Dove et al., 2022) and street flooding (Agonafir et al., 2022).
 Our investigation uses a 311 SR dataset covering 2-years (2022-2023), and was collected on 15 September 2024.
@@
 3 Data Cleansing Issues
-Data cleansing refers to the process of identifying and rectifying errors, inconsistencies, and inaccuracies within datasets to ensure they are of high quality and reliable for analysis (Maletic and Marcus,, 2005; Hosseinzadeh et al.,, 2023).
-The process typically involves removing duplicate records, handling missing or incomplete data, correcting mislabeled or inaccurate entries, and standardizing data formats (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018).
+Data cleansing refers to the process of identifying and rectifying errors, inconsistencies, and inaccuracies within datasets to ensure they are of high quality and reliable for analysis (Maletic and Marcus, 2005; Hosseinzadeh et al., 2023).
+The process typically involves removing duplicate records, handling missing or incomplete data, correcting mislabeled or inaccurate entries, and standardizing data formats (e.g., Cody, 2017; Van der Loo and De Jonge, 2018).
 In the context of open data, cleansing is especially important as open datasets often come from diverse, often uncoordinated sources, leading to variations in data quality, completeness, and consistency.
@@
 (a) Distribution of SR durations.
-(b) Post-closed resolution_action_update_date(s) ¿30 days
+(b) Post-closed resolution_action_update_date(s) >30 days
 Figure 4: Violin plots for negative SR durations and post-closed resolution updates (2022–2023.
@@
 As such, it offers dramatic improvements in file storage size, as well as efficient analysis and cross-platform support for R, Python, and Julia programming languages.
-Remarkable storage size reduction (within several orders of magnitude) is possible (Bates and Yan,, 2024).
+Remarkable storage size reduction (within several orders of magnitude) is possible (Bates and Yan, 2024).
 3.7 Data Dictionary
@@
 Trustworthy machine learning systems, which rely heavily on high quality data, are especially vulnerable to such issues.
-Errors in the underlying data, such as biases or inconsistencies, can compromise both the accuracy and fairness of predictive models, impacting real-world decisions in urban planning and policy (Rahm and Do,, 2000; Geiger et al.,, 2020).
+Errors in the underlying data, such as biases or inconsistencies, can compromise both the accuracy and fairness of predictive models, impacting real-world decisions in urban planning and policy (Rahm and Do, 2000; Geiger et al., 2020).
 While the findings here are specific to NYC 311 data, the principles proposed have broader applicability.
@@
 This challenge is even more pronounced in datasets that span long time periods, as changes in agency structures or reporting standards which can introduce subtle biases and data distortions.
-Long-term data consistency is crucial for longitudinal studies and predictive modeling, as even minor data shifts can lead to significant deviations in model outcomes, as noted by Rahm and Do, (2000) andBorgman, (2012).
+Long-term data consistency is crucial for longitudinal studies and predictive modeling, as even minor data shifts can lead to significant deviations in model outcomes, as noted by Rahm and Do (2000) andBorgman (2012).
 Finally, the intersection of data curation and machine learning for public policy applications opens new avenues for improving governance.
@@
 The COVID-19 pandemic underscored the importance of real-time, trustworthy data in crisis management, where the accuracy and timeliness of data-driven insights were critical for public health responses.
-Without proper data curation, machine learning models used during the pandemic would have been compromised, affecting decisions on resource allocation and service provision (Worby and Chang,, 2020; Khemasuwan and Colt,, 2021).
-Future efforts could focus on automating data curation processes, particularly in real-time data pipelines, to ensure that the data used in machine learning models remains accurate, clean, and reliable (Chu et al.,, 2016; Hurbean et al.,, 2021).
+Without proper data curation, machine learning models used during the pandemic would have been compromised, affecting decisions on resource allocation and service provision (Worby and Chang, 2020; Khemasuwan and Colt, 2021).
+Future efforts could focus on automating data curation processes, particularly in real-time data pipelines, to ensure that the data used in machine learning models remains accurate, clean, and reliable (Chu et al., 2016; Hurbean et al., 2021).
 Supplementary Material